# Difference-in-Discontinuities: Estimation, Inference and Validity Tests

### v1

@@ 2.2.1 Relaxing the Independence Assumption
$E\left[Y_{i,t}(1,1)-Y_{i,t}(1,0)-Y_{i,t}(0,1)+Y_{i,t}(0,0)|Z_{i}=z_{0}\right]\left(D_{0}^{+}D_{1}^{+}-D_{0}^{-}D_{1}^{-}\right)$
$\quad=E\left[Y_{i,t}(0,1)-Y_{i,t}(0,0)|Z_{i}=z_{0}\right]\left(D_{1}^{+}-D_{1}^{-}\right)E\left[Y_{i,t}(1,0)-Y_{i,t}(0,0)|Z_{i}=z_{0}\right]\left(D_{0}^{+}-D_{0}^{-}\right)$
This assumption allows for the possibility that the combined effect of the confounding policy and the treatment of interest is not simply the sum of their individual effects but it also includes an additional component that reflects how both treatments might interact. This additional component is determined by multiplying the effects together. By incorporating this concept, we can derive a new estimand that uses both the RDD at time 0 and the DiDC to identify the causal effect of the treatment of interest:
$\frac{\tau^{DiDC}}{1+\tau_{0}^{RDD}}=E\left[Y_{i,t}(0,1)-Y_{i,t}(0,0)|Z_{i}=z_{0}\right]\left(D_{1}^{+}-D_{1}^{-}\right)$
@@ 3 Estimation and Inference for the Sharp DiDC
Following standard practice in the regression discontinuity literature, we adopt local polynomial estimation to recover the parameter of interest. This nonparametric method involves fitting a polynomial function to the data near the threshold and using the estimated function to calculate the differences in outcomes between the treatment and control groups at the threshold.
To implement the estimation at $z_{0}$, we lean on the methodologies proposed by Calonico et al. (2014) for local polynomial estimation of the RDD estimator. In our case, we estimate the local polynomial regression on the differences in the outcomes with respect to time ($\Delta Y_{i}$).
3.1 Notation
Given a known threshold value $z_{0}$, which can be set to $z_{0}=0$ without loss of generality, the observed value $Z_{i}$ determines treatment assignment, with individuals assigned to the treatment group if $Z_{i}\geq 0$ and to the control group if $Z_{i}<0$ in the sharp setting.
The random variables $Y_{i,t}(1)$ and $Y_{i,t}(0)$ denote the potential outcomes of unit $i$ with and without treatment, respectively, at period $t$. The parameter of interest is $E\left[Y_{i}\left(1\right)-Y_{i}\left(0\right)|Z_{i}=z_{0}\right]$, the average treatment effect of treatment $D_{1}$ at the threshold. Under a mild continuity condition, Hahn et al. (2001) showed that $\tau^{RDD}=E[Y_{i}(1)-Y_{i}(0)|Z=z_{0}]$ is nonparametrically identifiable as the difference of two conditional expectations evaluated at the (induced) boundary point $z_{0}=0$. Similarly, the sharp DiDC parameter can be identified as the difference of the difference (in time) of two conditional expectations evaluated at $z_{0}=0$ at each side of $z_{0}$:
@@ 3.1 Notation
$\Delta\mu(z)$ $=E\left[\Delta Y_{i}|Z_{i}=z\right]$
We employ Assumptions A.1 and A.2 in Appendix A for the non-parametric local polynomial regression estimation. These assumptions impose restrictions on the kernel function as well as impose the existence of moments, enforce continuity of the running variable in the region, impose smoothness conditions on the regression functions, and restrict the conditional variance of the observed outcome.
@@ 3.2 Local Polynomial Estimator
For a given $\nu\leq p\in\mathbf{N}$, the general estimand of interest is $\tau^{DIDC}_{\nu}=\Delta\mu_{+}-\Delta\mu_{-}$ with $\Delta\mu^{(\nu)}_{+}=\nu!e^{\prime}_{\nu}\delta_{+,p}$, $\Delta\mu^{(\nu)}_{-}=\nu!e^{\prime}_{\nu}\delta_{-,p}$ being the $\nu$th-order derivatives of the $p$th-order local polynomial of the difference. The $p$th-order local polynomial estimators of the $\nu$th-order derivatives $\Delta\mu^{(\nu)}_{+,p}$ and $\Delta\mu^{(\nu)}_{-,p}$ are:
@@
$\hat{\tau}^{DiDC}_{\nu,p}(h_{n})=\Delta\hat{\mu}^{(\nu)}_{+,p}(h_{n})-\Delta\hat{\mu}^{(\nu)}_{-,p}(h_{n})$ (5)
@@ 3.2.1 Bandwidth choice
In order to perform the local polynomial estimation, it is necessary to choose an appropriate bandwidth $h_{n}$. This parameter determines the range of observations used for the estimation and impacts the trade-off between bias and variance in the estimated treatment effect $(\tau^{DiDC})$. When performing point estimation, the standard approach is to find the bandwidth that minimizes the asymptotic Mean Squared Error (MSE) of the estimator. We define the MSE as:
$MSE_{\nu,p,s}(h_{n})=E\left[\left(\hat{\tau}_{\nu,p,s}(h_{n})-\tau_{\nu,p}\right)^{2}|\chi_{n}\right]$
@@
Under Assumptions 1 and 2 with $S\geq p+1$, $\nu\leq p$, $h_{n}\rightarrow 0$ and $nh_{n}\rightarrow\infty$, the asymptotic MSE-optimal bandwidth is given by:
$h_{n}^{MSE}{}_{\nu,p}=\left(\frac{\left(1+2\nu\right)\mathbf{V}_{\nu,p}}{n2\left(1+p-\nu\right)\textbf{B}^{2}_{\nu,p,p+1,s}}\right)^{\frac{1}{2p+3}}$
@@ 3.3 Inference
$E\left[\hat{\tau}^{DiDC}_{\nu,p}(h_{n})|\chi_{n}\right]-\tau_{\nu}$ $=h_{n}^{p+1-\nu}\mathbf{B}_{\nu,p,p+1}(h_{n})+h_{n}^{p+2-\nu}\mathbf{B}_{\nu,p,p+2}(h_{n})$ (6)
@@
[…]
@@
$B\left[\hat{\tau}^{DiDC}_{\nu,p}(h_{n})\right]$ $=B\left[\hat{\tau}^{RDD}_{1,\nu,p}(h_{n})\right]-B\left[\hat{\tau}^{RDD}_{0,\nu,p}(h_{n})\right]$ (7)
where $B\left[\hat{\tau}^{RDD}_{1,\nu,p}(h_{n})\right]$ is the bias of the RDD estimated at time $t=1$, after the intervention happened, and $B\left[\hat{\tau}^{RDD}_{0,\nu,p}(h_{n})\right]$ is the bias of the RDD estimated at time $t=0$, before the intervention happened.
@@
Equation 7 demonstrates how incorporating additional data can help reduce the bias in RDD analysis. Traditional RDD analysis of interventions typically uses only solely on a cross-section of post-intervention data. However, by incorporating pre-intervention data into the analysis, bias can be substantially reduced and, under specific conditions, eliminated. Even in cases where the bias is not completely eliminated, it can still be improved if the RDD estimation of pre-treatment data exhibits bias in the same direction as that of post-treatment RDD and if its magnitude is not too large to overcome the original bias.
@@ 3.3.1 Bias Correction
[…]
@@
$b_{n}=h_{n}^{MSE}{}_{p+1,q,\nu+p+1}=\left(\frac{\left(3+2p\right)\mathbf{V}_{p+1,q}}{n2\left(q-p\right)\textbf{B}^{2}_{p+1,q,q+1,\nu+p+1}}\right)^{\frac{1}{2q+3}}$
3.3.2 Robust Confidence Interval
Following Calonico et al. (2014), we derive a large-sample distributional approximation that accounts for the added variability introduced by the bias estimate. The large-sample approximation for the standardized t-statistic is, if $S\geq q+1$, $n\min\{h_{n}^{2p+3},b_{n}^{2p+3}\}\times\max\{h_{n}^{2},b_{n}^{2(q-p)}\}\rightarrow\infty$, then
[…]
where $\mathbf{V}^{bc}_{\nu,p,q}\left(h_{n},b_{n}\right)$ is described in the appendix C. This justifies the following confidence intervals:
@@ 3.3.2 Robust Confidence Interval
These confidence intervals have better properties compared to conventional bias-corrected intervals. They are more robust to bandwidth selection, feature coverage error decays at a faster rate, and offer shorter interval lengths, as explained by Calonico et al. (2014).
4 Validity Test
@@ 4.1 Testing if Confounding Effect is Time-Invariant
As highlighted in Section 2, the importance of Assumption 3 - that the confounding effect is constant over time - cannot be overstated. Without this assumption, all estimations become inherently flawed, resulting in biased treatment effect estimates. Therefore, an initial step when considering the applicability of a differences-in-discontinuity approach is to assess the validity of this assumption.
To accomplish this, we propose a simple test to check violations of this assumption. In essence, it involves using the available pre-treatment periods to conduct an estimation of stacked RDDs. The goal is to examine whether the RDD coefficients remain consistent across multiple periods. This involves a set of $k$ periods where it is known that no changes occurred at the threshold.
@@
The procedure involves considering the following stacked RDDs regression model:
where $T_{-k}$ is a dummy indicating the period to which the data belong, that is, the RDD of which period. The hypothesis to be tested is:
@@
A question arises about concerning bandwidth to use when conducting this test. There are several options available, and Appendix D.1 provides details on the simulations used to assess the better bandwidth. In summary, in practice, any bandwidth derived from the data that is optimal for the specific RDD should work well.
4.2 Testing the Time-Invariance of Data-Generating Functional Form
In Section 3.3, we briefly discussed an important aspect of the DiDC method: its potential to achieve zero asymptotic bias when choosing DiDC over the standard RDD, provided that the functional forms of the data-generating processes for both groups are time-invariant. Essentially, this would mean that the derivatives of the functions on each side of the threshold at each $t=\{0,1\}$ ($\mu_{+,1}(z)$ and $\mu_{+,0}(z)$, $\mu_{-,1}(z)$ and $\mu_{-,0}(z)$) are equal at each point of the running variable $z$, resulting in zero bias in the local polynomial estimation.
Claim 3.
If the data-generating processes for both groups (above and below the threshold) are time-invariant, then $\Delta\mu_{+}^{(\nu)}$ and $\Delta\mu_{-}^{(\nu)}$ will be zero.

### v2

@@ 2.2.1 Relaxing the No-Interaction Assumption
$\mathbb{E}\left[Y_{i,1}(1,1)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]$
$\quad=\mathbb{E}\left[Y_{i,1}(0,1)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]\mathbb{E}\left[Y_{i,1}(1,0)-Y_{i,1}(0,0)|Z_{i}=z_{0}\right]$
This assumption allows for the possibility that the combined effect of the confounding policy and the treatment of interest is not simply the sum of their individual effects, but also includes an additional component reflecting their interaction. This additional component is determined by multiplying the effects together. By incorporating this concept, we can derive a new estimand that uses both the RDD at time 0 and the DiDC to identify the causal effect of the treatment of interest.
Lemma 3.
Under Assumptions 1, 2 of the policy of interest at the cutoff can be identified as3, 4 and 4’, the effect of the policy of interest at the cutoff can be identified as
$\frac{\tau^{DiDC}}{\tau^{RD}_{0}}+1\;=\;\mathbb{E}\!\left[Y_{i,1}(0,1)-Y_{i,1}(0,0)\;\middle|\;Z_{i}=z_{0}\right].$
Proof.
See Appendix C. ∎
@@ 3 Estimation and Inference for the Sharp DiDC
Following standard practice in the regression discontinuity literature, we propose a local polynomial estimation to recover the parameter of interest. This nonparametric method involves fitting a polynomial to the data near the threshold and using the estimated function to calculate differences in outcomes between the treatment and control groups at that threshold.
To implement the estimation at $z_{0}$, we rely on the methodologies proposed by Calonico et al. (2014b) for local polynomial estimation of the RDD. In our case, we estimate a local polynomial regression of the differences in outcomes over time ($\Delta Y_{i}$).
Under a mild continuity condition, Hahn et al. (2001) showed that the average treatment effect at the threshold is nonparametrically identifiable as the difference of two conditional expectations evaluated at the (induced) boundary point $z_{0}=0$. Similarly, the sharp DiDC parameter can be identified as the difference of the difference (in time) of two conditional expectations evaluated at $z_{0}=0$ at each side of $z_{0}$:
@@
$\Delta\mu(z)$ $=\mathbb{E}\left[\Delta Y_{i}|Z_{i}=z\right]$
Appendix A states the assumptions underlying the nonparametric local polynomial regression estimation. These assumptions impose restrictions on the kernel function, require the existence of certain moments, ensure continuity of the running variable in the relevant region, impose smoothness conditions on the regression functions, and bound the conditional variance of the observed outcome.
@@ 3.1 Local Polynomial Estimator
[…]
$\Delta\hat{\mu}^{(\nu)}_{+,p}(h_{n})$ $=\nu!e^{\prime}_{\nu}\hat{\delta}_{\Delta Y+,p}(h_{n})$
$\Delta\hat{\mu}^{(\nu)}_{-,p}(h_{n})$ $=\nu!e^{\prime}_{\nu}\hat{\delta}_{DeltaY-,p}(h_{n})$
@@
$\hat{\tau}^{DiDC}_{p}(h_{n})=\Delta\hat{\mu}_{+,p}(h_{n})-\Delta\hat{\mu}_{-,p}(h_{n})$ (4)
@@ 3.1.1 Bandwidth choice
To perform local polynomial estimation, it is necessary to choose an appropriate bandwidth $h_{n}$. This parameter determines the range of observations used for the estimation and impacts the trade-off between bias and variance in the estimated treatment effect $(\tau^{DiDC})$. In point estimation, the standard approach is to select the bandwidth that minimizes the asymptotic Mean Squared Error (MSE) of the estimator. Let $\chi_{n}=\begin{bmatrix}Z_{1}&\ldots&Z_{n}\end{bmatrix}^{\prime}$, the MSE is:
$MSE_{\nu,p,s}(h_{n})=\mathbb{E}\left[\left(\hat{\tau}_{\nu,p,s}(h_{n})-\tau_{\nu,p}\right)^{2}|\chi_{n}\right]$
@@
Under Assumptions 1 and 3 with $S\geq p+1$, $\nu\leq p$, $h_{n}\rightarrow 0$ and $nh_{n}\rightarrow\infty$, the asymptotic MSE-optimal bandwidth is given by:
$h_{n,\nu,p}^{MSE}=\left(\frac{\left(1+2\nu\right)\mathbf{V}_{\nu,p}}{2n\left(1+p-\nu\right)\textbf{B}^{2}_{\nu,p,p+1,s}}\right)^{\frac{1}{2p+3}}$
@@ 3.2 Inference
$\mathbb{E}\left[\hat{\tau}^{DiDC}_{\nu,p}(h_{n})|\chi_{n}\right]-\tau_{\nu}$ $=h_{n}^{p+1-\nu}\mathbf{B}_{\nu,p,p+1}(h_{n})+h_{n}^{p+2-\nu}\mathbf{B}_{\nu,p,p+2}(h_{n})$ (5)
@@
This is where we believe one of the main contributions of this research lies: the asymptotic bias of the DiDC can be zero if we include an assumption similar to that of parallel trends, or if the shapes of the data-generating processes for both groups are time-invariant. In the first case, imposition of “parallel trends“ restricts the functional form in such a way that $\Delta\mu_{+}^{(r)}\mathcal{B}_{+,\nu,p,r}(h_{n})=\Delta\mu_{-}^{(r)}\mathcal{B}_{-,\nu,p,r}(h_{n})$. It’s important to note that the symmetry of the kernel function, as imposed in Assumption A.2, plays a significant role in this result. In the case of time-invariant data-generating processes, both $\Delta\mu_{+}^{(\nu)}$ and $\Delta\mu_{-}^{(\nu)}$ equate to zero.
@@
$B\left[\hat{\tau}^{DiDC}_{\nu,p}(h_{n})\right]$ $=B\left[\hat{\tau}^{RD}_{1,\nu,p}(h_{n})\right]-B\left[\hat{\tau}^{RD}_{0,\nu,p}(h_{n})\right]$ (6)
where $B\left[\hat{\tau}^{RD}_{1,\nu,p}(h_{n})\right]$ is the bias of the RD estimated at time $t=1$, after the intervention happened, and $B\left[\hat{\tau}^{RD}_{0,\nu,p}(h_{n})\right]$ is the bias of the RD estimated at time $t=0$, before the intervention happened.
@@
Equation 6 demonstrates how incorporating additional data can help reduce the bias in RD analysis. Traditional RD analysis of interventions typically uses only a cross-section of post-intervention data. However, by incorporating pre-intervention data into the analysis, bias can be substantially reduced and, under specific conditions, eliminated. Even when the bias is not completely eliminated, it can still be reduced if the RD estimation of pre-treatment data exhibits bias in the same direction as the post-treatment RD and if its magnitude is not too large to overcome the original bias.
@@ 3.2.1 Bias Correction
Using the MSE-optimal bandwidth for point inference can result in bandwidths that are “too large“, which may seem attractive for reducing variance but can introduce first-order asymptotic bias [^1] Calonico et al. (2014b). To address this issue, we employ robust bias-corrected confidence intervals (CIs) proposed by Calonico et al. (2014b). These intervals take into account the asymptotic bias of the point estimate by 1) estimating the bias and recentering the CI, and 2) incorporating the additional variance from estimating the bias for bias correction into the CI. This process requires estimating a separate local polynomial of order $q$, with $q>p\geq\nu$. The bias-corrected estimator is defined as:
@@
$b_{n}=h_{n,p+1,q,\nu+p+1}^{MSE}=\left(\frac{\left(3+2p\right)\mathbf{V}_{p+1,q}}{2n\left(q-p\right)\textbf{B}^{2}_{p+1,q,q+1,\nu+p+1}}\right)^{\frac{1}{2q+3}}$
3.2.2 Asymptotic Properties and Robust Confidence Interval
Following Calonico et al. (2014b), we derive a large-sample distributional approximation that accounts for the added variability introduced by the bias estimate. The large-sample approximation for the standardized t-statistic is formalized in the theorem below:
Theorem 1.
Under Assumptions 1-5, if $S\geq q+1$, $n\min\{h_{n}^{2p+3},b_{n}^{2p+3}\}\times\max\{h_{n}^{2},b_{n}^{2(q-p)}\}\rightarrow\infty$, then
$T^{rbc}_{\nu,p,q}\left(h_{n},b_{n}\right)=\frac{\hat{\tau}^{bc}_{\nu,p,q}\left(h_{n},b_{n}\right)-\tau_{\nu}}{\sqrt{\mathbf{V}^{bc}_{\nu,p,q}\left(h_{n},b_{n}\right)}}\xrightarrow{d}\mathcal{N}\left(0,1\right)$
where $\mathbf{V}^{bc}_{\nu,p,q}\left(h_{n},b_{n}\right)$ is described in appendix C.
This motivates the following confidence interval:
@@ 3.2.2 Asymptotic Properties and Robust Confidence Interval
Just like in the standard RD setting, this confidence interval have better properties compared to conventional bias-corrected intervals. They are more robust to bandwidth selection, feature coverage error decays at a faster rate, and offer shorter interval lengths, as explained by Calonico et al. (2014b).
4 Validity Tests
@@ 4.1 Testing if Confounding Effect is Time-Invariant
As highlighted in Section 2, the importance of Assumption 4 - that the confounding effect is constant over time - cannot be overstated. Without this assumption, all estimations are inherently flawed, leading to biased treatment effect estimates. Therefore, an initial step when considering the applicability of a differences-in-discontinuity approach is to assess the validity of this assumption.
To accomplish this, we propose a simple test to check violations of this assumption. In essence, it involves using available pre-treatment periods to estimate stacked RDs. The goal is to examine whether the RD coefficients remain consistent across multiple periods. This involves a set of $k$ periods where it is known that no changes occurred at the threshold.
@@
The procedure involves considering the following stacked RDs regression model:
$Y_{i}=\alpha_{i}+\sum_{k=1}^{K}\left[\beta_{-k}\left(Z_{i}-z_{0}\right)+\theta_{-k}D_{i}\left(Z_{i}-z_{0}\right)\right]T_{-k}+\varepsilon_{i}$ (7)
where $T_{-k}$ is a dummy indicating the period to which the data belong, that is, the RD of which period. The hypothesis to be tested is:
@@
A question arises about concerning bandwidth to use when conducting this test. There are several options available, and Appendix D.1 provides details on the simulations used to assess the better bandwidth. In summary, in practice, any bandwidth derived from the data that is optimal for the specific RD should work well.
4.2 Testing the Time-Invariance of Potential Outcomes Functional Form
In Section 3.2, we briefly discussed an important aspect of the DiDC method: its potential to achieve zero asymptotic bias when choosing DiDC over the standard RD, provided that the functional forms of the data-generating processes for both groups are time-invariant. Essentially, this would mean that the derivatives of the functions on each side of the threshold at each $t=\{0,1\}$ ($\mu_{+,1}(z)$ and $\mu_{+,0}(z)$, $\mu_{-,1}(z)$ and $\mu_{-,0}(z)$) are equal at each point of the running variable $z$, resulting in zero bias in the local polynomial estimation.